OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1

OpenStudy (anonymous):

(3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (anonymous):

@shinalcantara

OpenStudy (igreen):

-2x + y = 1 Add 2x to both sides: y = 2x + 1 We can plug in 2x + 1 for 'y' in the 2nd equation: -4x + 2x + 1 = -1 Add: -2x + 1 = -1 Subtract 1 to both sides: -2x = -2 Divide -2 to both sides. What's -2 / -2? @domisreallytall

OpenStudy (igreen):

@domisreallytall Can you divide -2 and -2?

OpenStudy (anonymous):

-1?

OpenStudy (igreen):

No, not quite.. -2 / -2 is 1..so that's our y-value.

OpenStudy (anonymous):

so its a?

OpenStudy (igreen):

Sorry, 1 is our x-value..

OpenStudy (igreen):

Not y..

OpenStudy (anonymous):

oh. so its D

OpenStudy (igreen):

Yep, you got it.

OpenStudy (anonymous):

Thanks can you help me with 3 more?